# Exploring the Adversarial Vulnerabilities of Vision-Language-Action Models in Robotics

### v2

@@ 4.3 Diagnostic Experiment
Impact of Patch Size. We further study TMA’s performance with different patch sizes, and report the average failure rate across four LIBERO [46] benchmark tasks in Fig. 4(b). The patch sizes examined were $[3\%,5\%,7\%,10\%]$ of a $224\times 224$ input image. Our findings indicate an inversely proportional relationship between the $L_{1}$ distance of predicted actions and the target action as the patch size increased. This increase in patch size also correlated with a marked rise in task failure rates. The observed trend suggests that larger patch sizes afford the adversary more extensive optimization space to influence the model’s output, which is consistent with previous work [12, 84].
@@ 4.4 Robustness Evaluation
We further examine if concurrent defense strategies [18, 83, 81] can generalize against our adversarial examples introduced within VLA. Specifically, we applied four prior defense techniques, and reported their task failure rates for both benign and adversarial samples in Fig. 4, respectively. The results indicate that our adversarial attack can escape from most of the defense strategies. More details of defense parameters see in §S7.
@@ 4.5 Systemic Discussion
Figure 6: Patch Visualization. Sematic-rich patches generated from physical or simulation datasets.
@@ 5 Conclusion
While VLA models have gained significant popularity due to their substantial robot capabilities, this study, for the first time, establishes specific robotic attack objectives and demonstrates that these models are in fact vulnerable to various types of attacks. Experimental results demonstrate that our attacks expose significant vulnerabilities to VLA models, raising concerns regarding impetuous real-world deployments. We believe our framework provides pioneering and foundational contributions to the reliability of AI enhanced robotic systems.
6 Acknowledgements
This research was supported by the National Science Foundation under Grant No. 2242243. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of U.S. Naval Research Laboratory (NRL) or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
Appendix S1 Black-box Attack Results
@@ 6 Acknowledgements
Appendix S2 Targeted Attack supply Results
In this section, we explore Targeted Manipulation Attack (TMA) with different action magnitude targets and investigate the impact of patch size on failure rate.
S2.1 Investigation of Action Magnitude
@@
S2.2 Investigation of Patch Size
Acknowledging the trade-off between stealth and attack effectiveness [8, 84, 45], in this section, we further examine the impact of different patch sizes on AFR to explore how this balance influences performance.
Table S3 presents the average failure rates of the TMA attack on DoF1 with a target action magnitude of 0 under the physical setup. The results indicate that the failure rate increases as the patch size grows. Balancing stealth and attack efficiency, we adopt a patch size of 5% as the primary attack configuration for all experiments.
Table S3: Patch Size. We compare the average failure rates across four LIBERO [46] tasks using different patch sizes. Patch size (%) indicates the proportion of the adversarial patch to the input image.
Setup Patch Size (%) AFR (%) Physical 3 72.4 5 89.1 7 92.4 10 93.6
@@
Appendix S3 Details of Real-world Experiments
@@
Appendix S4 Failure Case Analysis
@@
Appendix S5 More Evaluation Details
This section supplies details on the evaluation of the three attacks in §4.1. To rigorously assess the effectiveness of attacks and minimize task failures arising from random placement of adversarial patches that obscure critical task-relevant objects, we identify base left-corner locations for patch placement specific to each of the four tasks in LIBERO [46]. These locations are further extended by shifting the x and y coordinates by $\pm$3 pixels, resulting in a total of 9 placement locations. For the Spatial task, the base point is set as $(120,160)$. For the Object task, the base point is $(30,150)$. Similarly, the base point is $(15,158)$ for the Goal task and $(5,160)$ for the Long task. This approach ensures an accurate evaluation of adversarial patches on task performance.
Appendix S6 More Diagnostic Experiment Details
@@
Appendix S7 More Robustness Evaluation Details
Appendix S8 More Qualitative Results
This section presents additional qualitative results for each of the three attack objectives, complementing results in §4.2 and offering deeper insights into the effectiveness of our adversarial attacks. Specifically, the results of UADA, UPA, and TMA are shown in Fig. S2S3, and S4, respectively. These qualitative results reveal significant deviations in the adversarial trajectories of the proposed methods across all four LIBERO tasks compared to the benign trajectories. Specifically, UADA induces substantial action discrepancies while diverging sharply from the benign trajectory, highlighting its disruptive impact. For UPA, a significant position deviation is observed on the spatial task [46], where the adversarial trajectory deviates significantly from the benign trajectory, ultimately causing task failure. In contrast, the TMA exhibits a smaller action amplitude, successfully manipulating the robot’s behavior and resulting in task failure. These findings highlight the distinct impacts and underlying mechanisms of the three attack methods.
Appendix S9 Future direction

### v3

@@ 4.4 Diagnostic Experiment
Impact of Patch Size. We further study TMA’s performance with different patch sizes and report the average FR across four LIBERO [44] benchmark tasks in Fig. 5. The patch sizes examined were $[3\%,5\%,7\%,10\%]$ of a $224\times 224$ input image. Our findings indicate an inversely proportional relationship between the $L_{1}$ distance of predicted actions and the target action as the patch size increased. This increase in patch size also correlated with a marked rise in FR. The observed trend suggests that larger patch sizes afford the adversary more optimization space to influence the model’s output, aligning with prior work [12, 81].
@@ 4.5 Robustness Evaluation
We further examine if concurrent defense strategies [18, 80, 78] can generalize against our adversarial examples introduced within VLA. Specifically, we applied four prior defense techniques and reported their FR for both benign and adversarial samples in Fig. 5, respectively. The results indicate that our adversarial attack can escape from most of the defense strategies. More details see in §S6.
@@ 4.6 Systemic Discussion
Figure 6: Patch Visualization. Sematic-rich patches.
@@ 5 Conclusion
While VLA models have gained significant popularity due to their substantial robot capabilities, this study, for the first time, establishes specific robotic attack objectives and demonstrates that these models are in fact vulnerable to various types of attacks. demonstrate that our attacks expose significant vulnerabilities to VLA models, raising concerns regarding impetuous real-world deployments. We believe our framework provides pioneering and foundational contributions to the reliability of AI-enhanced robotic systems.
Supplementary Material
SUMMARY OF THE APPENDIX
This appendix contains additional experimental results and discussions of our work, organized as:
  • §S1 includes additional related work in for robotic system safety.
  • §S2 includes additional simulation results for black-box attacks with different VLA-based models.
  • §S3 outlines the experiment details.
  • §S4 provides targeted attack supply results, investigating TMA across various action magnitudes and patch sizes.
  • §S5 explains tasks adopted for real-world experiments.
  • §S6 outlines the parameter configurations used in robustness evaluations.
  • §S7 studies failure cases, providing an in-depth analysis of the limitations of the proposed methods.
  • §S8 supplies additional information on diagnostic experiments.
  • §S9 visualizes more qualitative results.
  • §S10 provides discussions on future directions, highlighting potential areas for further research.
S1 Human-Robot Interaction Safety
S2 Black-box Attack Results
@@
S3 Experiment Details
Attacking Details: Following [34], we apply scaling and normalization for data preprocessing. The initial learning rate is set to $2e^{-3}$, with the AdamW optimizer [47] with a cosine annealing scheduler, including $20$ warm-up iterations. Training is conducted over $2e^{3}$ iterations with a batch size of $6$ and $50$ inner-loop steps (see Algorithm 1). The transformation parameters $\phi,\psi$ in Eq. 9 are set to $30$ and $0.2$ to simulate real-world perspective change. For all three objectives, considering the trade-off between stealth and performance observed in §S4, we set the patch size to $5\%$. For the hyperparameters $\alpha,\beta$ in Eq. 5, We conduct grid search and tune $\alpha$ from $[0.2,0.4,0.6,0.8]$ and $\beta$ from $[0.8,0.6,0.4,0.2]$. We select $\alpha=0.8$ and $\beta=0.2$ based on the validation set performance. Ablation studies $w.r.t.$ $\alpha$ and $\beta$ are presented in §S8.
Evaluation Details: To rigorously assess the effectiveness of attacks and minimize task failures arising from random placement of adversarial patches that obscure critical task-relevant objects, we identify base left-corner locations for patch placement specific to each of the four tasks in LIBERO [44]. For the Spatial task, the base point is set as $(120,160)$. For the Object task, the base point is $(30,150)$. Similarly, the base point is $(15,158)$ for the Goal task and $(5,160)$ for the Long task. Additionally, to enhance reproducibility by mitigating randomness, patches remain untransformed during evaluation. This approach ensures a reliable assessment of adversarial patches’ impact on task performance.
S4 Targeted Attack supply Results
@@
S5 Details of Real-world Experiments
@@
S6 More Robustness Evaluation Details
S7 Failure Case Analysis
@@
S8 More Diagnostic Experiment Details
@@
S9 More Qualitative Results
This section presents additional qualitative results for each of the three attack objectives, complementing results in §4.3 and offering deeper insights into the effectiveness of our adversarial attacks. Specifically, the results of UADA, UPA, and TMA are shown in Fig. S1S2, and S3, respectively. These qualitative results reveal significant deviations in the adversarial trajectories of the proposed methods across all four LIBERO tasks compared to the benign trajectories. Specifically, UADA induces substantial action discrepancies while diverging sharply from the benign trajectory, highlighting its disruptive impact. For UPA, a significant position deviation is observed on the spatial task [44], where the adversarial trajectory deviates significantly from the benign trajectory, ultimately causing task failure. In contrast, the TMA exhibits a smaller action amplitude, successfully manipulating the robot’s behavior and resulting in task failure. These findings highlight the distinct impacts and underlying mechanisms of the three attack methods.
S10 Future direction